# Perception, Reason, Think, and Plan: A Survey on Large Multimodal Reasoning Models

### v1

@@ 1 Introduction
  • This paper presents a comprehensive survey of the Large Multimodal Reasoning Model (LMRM) landscape, encompassing over 540 publications. Our analysis contextualizes and addresses key reasoning limitations in current models (Sec. 2).
@@
  • We reorganize existing datasets and benchmarks (update to 2025.04) of multimodal understanding and reasoning (Sec. 5) to clarify their categories and evaluation dimensions.
@@ 2 Evolving Paradigms of Multimodal Reasoning and Discussion
However, this unification primarily emphasized visual representation and cross-modal fusion, often at the expense of deeper semantic modelling of language. As a result, the reasoning process frequently defaulted to a classification-based paradigm, limiting context-aware and generalized reasoning. The multimodal reasoning systems still rely on additional modules or task-specific enhancements. Overall, the reasoning in this stage remained largely implicit by foundational perceptual processing and neural computation. The emerging multimodal language models will enhance implicit reasoning by introducing powerful language models and large-scale visual data.
@@
To address these gaps, we prospect the development of␣native large multimodal reasoning models (N-LMRMs) as a potential paradigm shift in machine intelligence4). In contrast to conventional LMRMs, which retrofit language models with auxiliary modality processors, N-LMRMs will be natively designed to unify multimodal understanding, generation, and agentic reasoning within a fully end-to-end architecture. Real-world data types are encoded within a unified representation space, like VideoPoet (Kondratyuk et al., 2024), while large-scale synthetic data facilitates holistic learning of reasoning and planning in the environment of any modality interaction. This evolution hinges on two transformative capabilities: 1) Multimodal Agentic Reasoning: N-LMRMs will embody agentic intelligence, enabling proactive, goal-driven interactions with complex environments, such as long-horizon planning—hierarchical task breakdown and memory-enhanced reasoning for coherence in extended interactions; dynamic adaptation—real-time strategy adjustment based on environmental feedback; embodied learning—closed-loop training frameworks enabling models to learn through simulated or physical interactions for better generalization. 2)␣Omni-Modal Understanding and Generative Reasoning: N-LMRMs will move beyond modality-specific encoders and decoders by utilizing a unified representational space for smooth cross-modal synthesis and analysis. This approach includes heterogeneous data fusion for the joint embedding of diverse data types, contextual multimodal generation for the coherent creation of composite outputs, and modality-agnostic inference that enables adaptable processing pipelines for the task-agnostic handling of new or any cross-modal data.
@@ 3.1 Stage 1 Perception Driven Modular Reasoning - Developing Task-Specific Reasoning Modules
BLIP-2 (Li et al., 2023d); LMEye (Li et al., 2024l) LISA (Lai et al., 2024);
@@
AGoT (Yang et al., 2024c); MM-ToT (Gomez, 2023) ,leaf, text width=26.5em, draw=c9, line width=0.7mm, edge=c9 ] ] [ Tools, fill=c9!60, draw=c9, line width=0.5mm, edge=c9 [ L3GO (Yamada et al., 2024); HYDRA (Ke et al., 2024);
@@
, leaf, text width=38em, draw=c4, line width=0.7mm, edge=c5 ] ] [ Multimodal-R1, fill=c4!60, draw=c4, line width=0.5mm, edge=c4 [ Deepseek-R1 (DeepSeek-AI et al., 2025); R1-OneVision (Yang et al., 2025e)
Rlhf-v (Yu et al., 2024a); Vt-r1 (Zhou et al., 2025); MMPR (Wang et al., 2024g);
@@
ManipLLM (Li et al., 2024i); RAGEN (Wang et al., 2025i)
@@
SeeClick (Cheng et al., 2024b); InfiGUI-R1 (Liu et al., 2025d)
@@
Kimi-VL-Thinking (Team, 2025a); Uni-MoE (Li et al., 2025j); AnyGPT (Zhan et al., 2024);
@@
GPT-4.1 (OpenAI, 2025a); M2-Omni (Guo et al., 2025); R1-Omni (Zhao et al., 2025c);
Baichuan-Omni-1.5 (Li et al., 2025h); Megrez-3B-Omni (Li et al., 2025a)
@@ 3.1.2 Vision-Language Models-based Modular Reasoning
BLIP-2 (Li et al., 2023d) 2023 Single-Transformer-Backbone Uses a querying Transformer for improved cross-modal reasoning efficiency. Pretraining + fine-tuning
@@
Single-Transformer-Backbone Interactive Reasoning. This paradigm embeds visual and textual inputs in a single Transformer, enabling direct cross-modal reasoning through unified encoding method. VisualBERT (Li et al., 2019), UNITER (Chen et al., 2020), VL-BERT (Su et al., 2019) fuse visual-text inputs in a single Transformer to reason via joint contextual encoding or enhanced cross-modal pretraining. PixelBERT (Huang et al., 2020) employs a CNN and Transformer architecture to process pixels for fine-grained reasoning (e.g., NLVR2). UniVL (Luo et al., 2020) unifies video-language reasoning with a single Transformer for temporal cross-modal tasks (e.g., TVQA). Oscar (Li et al., 2020), VinVL (Zhang et al., 2021a) anchor reasoning with object tags or enhanced visual features in a unified Transformer, boosting semantic inference (e.g., VCR, GQA). ERNIE-ViL (Yu et al., 2021) integrates scene graph knowledge into a single Transformer, enhancing compositional reasoning through structured visual-language interactions. UniT (Hu & Singh, 2021) streamlines multimodal tasks with a shared self-attention backbone for unified reasoning. PaLI (Chen et al., 2022b) scales single-Transformer reasoning with a multilingual framework for cross-lingual inference (e.g., OK-VQA). Flamingo (Alayrac et al., 2022) employs cross-attention to prioritize dynamic vision-text interactions. BEiT-3 (Wang et al., 2022b) adopts masked data modeling to unify vision-language learning. OFA (Wang et al., 2022a), BLIP-2 (Li et al., 2023d) introduce a unified multimodal framework or a querying Transformer to excel in cross-modal reasoning with improved efficiency (e.g., VQA-v2). Kosmos-1 (Huang et al., 2023b), Kosmos-2 (Peng et al., 2023) enable interleaved input processing or grounding capability for flexible multimodal understanding and precise object localization.
Multimodal LLMs-based Implicit Reasoning. This approach projects visual inputs into a large language model’s text space, leveraging the contextual reasoning capabilities of large language models (Li et al., 2023e) to improve the performance of multimodal reasoning. Their architecture contains pretrained visual encoders and large language models, arr. Vision-Encoder-LLM. CLIP-Cap (Mokady et al., 2021) projects CLIP visual features into an LLM for reasoning and captioning tasks. LLaVA (Liu et al., 2023a) enables conversational reasoning by tuning ViT-LLM integration for interactive tasks or scaling for complex VQA tasks. MiniGPT-4 (Zhu et al., 2023), InstructBLIP (Dai et al., 2023) align a ViT to a frozen LLM via a projection layer or instruction tuning, streamlining visual-text reasoning. Qwen-VL (Bai et al., 2023) incorporates a spatial-aware ViT, enhancing grounded reasoning for spatially complex tasks. mPLUG-Owl (Ye et al., 2023), LMEye (Li et al., 2024l), and Otter (Li et al., 2023a) integrate a modular visual encoder with an LLM for instruction-following and in-context learning for multimodal reasoning.
@@ 3.2.1 Prompt-based MCoT
Prompt-based Multimodal Chain-of-Thought (MCoT) methods extend the textual CoT paradigm to multimodal contexts, enabling step-by-step reasoning across modalities with strong interpretability and minimal additional training. In visual reasoning, IPVR (Chen et al., 2023c) proposed a structured “see-think-confirm” prompting framework that guides LLMs through visual grounding and rationale verification. VIC (Zheng et al., 2024c) prompts textual reasoning chains before visual input to mitigate hallucinations and improve accuracy.
@@ 3.2.2 Structural Reasoning
Unlike prompt-based MCoT methods, which induce reasoning behaviour through handcrafted exemplars or a zero-shot prompting approach, structural reasoning focuses on learning reasoning patterns through supervised training. By integrating explicit procedural structures into the model, these approaches transform loosely guided reasoning into standardized, stage-wise processes, improving scalability, reliability, and efficiency in complex multimodal tasks. We categorize structural reasoning into three representative types: (i) rationale construction, which learns to produce atomic reasoning steps as interpretable scaffolds; (ii) defined reasoning procedures, which adapt structured texture reasoning schemes to multimodal settings; and (iii) modality-specific structural reasoning, which further incorporates modality-aware constraints and designs to better align with the characteristics of visual, auditory, or embodied inputs.
@@
The foundation of structural reasoning in multimodal contexts begins with effective rationale learning approaches. Multimodal-CoT (Zhang et al., 2023g) proposes a two-stage Multimodal-CoT framework that decouples rationale generation from answer prediction to reduce hallucinations. T-sciq (Wang et al., 2024e) leverages teacher LLMs to generate rationales with varying complexity, showing rationale quality is key to reasoning accuracy. In autonomous driving, G-CoT (Ma et al., 2024) designs Dolphins explicitly linking rationales to visual and historical driving signals for more grounded reasoning. MC-CoT (Tan et al., 2024a) uses a self-consistency strategy to select the most accurate rationale among multiple candidates, boosting smaller models’ performance. CLoT (Zhong et al., 2024a) promotes non-linear, explorative rationale construction via Leap-of-Thought to support creative reasoning.
@@ 3.2.3 Externally Augmented Reasoning
Externally augmented reasoning introduces advantage algorithm, auxiliary tools or expert modules to compensate for limitations in the model’s inherent reasoning capacity. These components are integrated at inference time or coupled during training, enabling more flexible, scalable, and task-specialized reasoning workflows. By decoupling core reasoning steps from the base model, such methods support long-horizon planning, precise grounding, and access to dynamic or domain-specific information. We group externally augmented methods into four categories: (i) search algorithm-enhanced MCoT, which navigates reasoning spaces via various search algorithm; (ii) tool-based augmentation, which leverages external language tools or systems to guide reasoning execution; (iii) retrieval-augmented reasoning, which incorporates relevant multimodal knowledge from external sources into the reasoning path; and (iv) multimodal enhancing, which incorporate specialized multimodal modules to support perception-driven reasoning.
@@
AGoT  (2024c) T,I Text-Image Retrieval, VQA Search Algorithm prompt aggregation and prompt flow operations Builds a graph to aggregate multi-faceted reasoning with visuals.
@@
Search strategy-driven MCoT approaches empower models to dynamically navigate and optimize reasoning trajectories throughout the reasoning process. MM-ToT (Gomez, 2023), for instance, leverages GPT-4 and Stable Diffusion, employing depth-first search (DFS) and breadth-first search (BFS) algorithms to identify the most optimal multimodal outputs according to a 0.0–1.0 metric scale. HoT (Yao et al., 2023a) creates interconnected thoughts from multimodal inputs and packages them into a single hyperedge. Unlike this, Aggregation Graph-of-Thought (AGoT) (Yang et al., 2024c) builds a reasoning aggregation graph, which integrates diverse reasoning elements at every step and subsequently incorporates visual data. Blueprint Debate on Graph (BDoG) (Zheng et al., 2024b) takes a distinctive route, discarding search algorithms and instead utilizing three agents—an affirmative debater, a negative debater, and a moderator. These agents engage in iterative debates to address multimodal questions, with the moderator ultimately synthesizing a final answer, thus implicitly constructing a graph-of-thought that explores and aggregates a wide range of thoughts. Overall, compared to prompt-based methods that rely on linear, example-driven inference, search strategy-oriented MCoT variants enable models to explore multiple reasoning pathways, thereby significantly enhancing adaptability and the depth of problem-solving.
@@ 3.3.1 Cross-Modal Reasoning
Cross-Modal Reasoning refers to the ability to integrate and reason across multiple modalities, such as text, images, videos. Recent advancements in cross-modal reasoning have emphasized the importance of augmenting multimodal information beyond textual inputs through model-intrinsic capabilities or external tools and algorithms. These methods aim to enhance reasoning accuracy and robustness by dynamically incorporating complementary information from diverse modalities.
@@
Beyond the use of external tools for multimodal understanding described in §3.2.3, recent approaches increasingly explore tool integration as a vehicle for multimodal reasoning itself. VisProg (Gupta & Kembhavi, 2023) and ProViQ (Choudhury et al., 2024) leverage program generation and procedural execution to enable cross-modal reasoning, dynamically generating executable code or logic paths to solve complex tasks such as video question answering, multi-step visual reasoning, and geometric problem solving. In parallel, methods such as AssistGPT (Gao et al., 2023), MM-ReAct (Yang et al., 2023), and Multi-Modal-Thought (Lin et al., 2025a) adopt modular integration frameworks—such as PEIL and vision expert prompting—to coordinate tool use based on reasoning progression. These systems enable interpretable and adaptive reasoning by calling different tools dynamically during task execution. VisualReasoner (Cheng et al., 2024a) further introduces a data synthesis strategy to generate multi-step reasoning traces, which are then used to train a plug-and-play visual reasoning module applicable to a variety of vision-language backbones. Collectively, these efforts extend the landscape of multimodal reasoning by combining program induction, dynamic tool orchestration, and data-driven reasoning supervision.
@@
FAST (Sun et al., 2024a) and ICoT (Gao et al., 2024a) both leverage cognitive processes akin to human thinking, with FAST employing a system switch adapter to dynamically alternate between fast and slow thinking modes, while ICoT utilizes Attention-driven Selection (ADS) to interleave visual and textual reasoning steps. Meanwhile, Image-of-Thought (Zhou et al., 2024b) and CoTDiffusion (Ni et al., 2024a) focus on generating visual rationales, with Image-of-Thought extracting visual info step-by-step and CoTDiffusion creating visual subgoal plans, extending algorithmic augmentation to robotics.
@@
These approaches rely on the LMM’s inherent ability to generate or infer multimodal information without external tools. T-SciQ (Wang et al., 2024e), Visual-CoT (Rose et al., 2023) and VoCoT (Li et al., 2024m) demonstrated how fine-tuning LMMs on carefully designed CoT datasets (e.g., VoCoT-Instruct80K) could enable single-step multimodal reasoning in charts, documents, and geometry problems. MVoT (Li et al., 2025b) represents an early effort, where a self-contained architecture iteratively refines visual-textual representations for embodied reasoning tasks.
@@
Table 5: Apporach of Multimodal-o1. It mainly relies on a multi-stage, structured reasoning path to solve problems.

### v2

@@ 1 Introduction
  • This paper presents a comprehensive survey of the Large Multimodal Reasoning Model (LMRM) landscape, encompassing about 700 publications. Our analysis contextualizes and addresses key reasoning limitations in current models (Sec. 2).
@@
  • We reorganize existing datasets and benchmarks (update to 2025.06.30) of multimodal understanding and reasoning (Sec. 5) to clarify their categories and evaluation dimensions.
@@ 2 Evolving Paradigms of Multimodal Reasoning and Discussion
However, this unification primarily emphasized visual representation and cross-modal fusion, often at the expense of deeper semantic modelling of language. As a result, the reasoning process frequently defaulted to a classification-based paradigm, limiting context-aware and generalized reasoning. The multimodal reasoning systems still rely on additional modules or task-specific enhancements. Overall, reasoning at this stage remained largely implicit, primarily driven by foundational perceptual processing and neural computation. The emerging multimodal language models will enhance implicit reasoning by introducing powerful language models and large-scale visual data.
@@
To address these gaps, we prospect the development of␣native large multimodal reasoning models (N-LMRMs) as a potential paradigm shift in machine intelligence4). In contrast to conventional LMRMs, which retrofit language models with auxiliary modality processors, N-LMRMs will be natively designed to unify multimodal understanding, generation, and agentic reasoning within a fully end-to-end architecture. Real-world data types are encoded within a unified representation space, like VideoPoet (Kondratyuk et al., 2024), BAGEL (Deng et al., 2025a), and Janus-Pro (Chen et al., 2025g), while large-scale synthetic data facilitates holistic learning of reasoning and planning in the environment of any modality interaction. This evolution hinges on two transformative capabilities: 1) Multimodal Agentic Reasoning: N-LMRMs will embody agentic intelligence, enabling proactive, goal-driven interactions with complex environments, such as long-horizon planning—hierarchical task breakdown and memory-enhanced reasoning for coherence in extended interactions; dynamic adaptation—real-time strategy adjustment based on environmental feedback; embodied learning—closed-loop training frameworks enabling models to learn through simulated or physical interactions for better generalization. 2)␣Omni-Modal Understanding and Generative Reasoning: N-LMRMs will move beyond modality-specific encoders and decoders by utilizing a unified representational space for smooth cross-modal synthesis and analysis. This approach includes heterogeneous data fusion for the joint embedding of diverse data types, contextual multimodal generation for the coherent creation of composite outputs, and modality-agnostic inference that enables adaptable processing pipelines for the task-agnostic handling of new or any cross-modal data.
@@ 3.1 Stage 1 Perception Driven Modular Reasoning - Developing Task-Specific Reasoning Modules
BLIP-2 (Li et al., 2023c); LMEye (Li et al., 2024l) LISA (Lai et al., 2024);
@@
AGoT (Yang et al., 2024d); MM-ToT (Gomez, 2023) ,leaf, text width=26.5em, draw=c9, line width=0.7mm, edge=c9 ] ] [ Tools, fill=c9!60, draw=c9, line width=0.5mm, edge=c9 [ L3GO (Yamada et al., 2024); HYDRA (Ke et al., 2024);
@@
, leaf, text width=38em, draw=c4, line width=0.7mm, edge=c5 ] ] [ Multimodal-R1, fill=c4!60, draw=c4, line width=0.5mm, edge=c4 [ Deepseek-R1 (DeepSeek-AI et al., 2025); R1-OneVision (Yang et al., 2025g)
Rlhf-v (Yu et al., 2024a); Vt-r1 (Zhou et al., 2025b); MMPR (Wang et al., 2024g);
@@
ManipLLM (Li et al., 2024i); RAGEN (Wang et al., 2025z)
@@
SeeClick (Cheng et al., 2024b); InfiGUI-R1 (Liu et al., 2025i)
@@
Kimi-VL-Thinking (Team, 2025a); Uni-MoE (Li et al., 2025n); AnyGPT (Zhan et al., 2024);
@@
GPT-4.1 (OpenAI, 2025a); M2-Omni (Guo et al., 2025); R1-Omni (Zhao et al., 2025d);
Baichuan-Omni-1.5 (Li et al., 2025k); Megrez-3B-Omni (Li et al., 2025a)
@@ 3.1.2 Vision-Language Models-based Modular Reasoning
BLIP-2 (Li et al., 2023c) 2023 Single-Transformer-Backbone Uses a querying Transformer for improved cross-modal reasoning efficiency. Pretraining + fine-tuning
@@
Single-Transformer-Backbone Interactive Reasoning. This paradigm embeds visual and textual inputs in a single Transformer, enabling direct cross-modal reasoning through unified encoding method. VisualBERT (Li et al., 2019), UNITER (Chen et al., 2020), VL-BERT (Su et al., 2019) fuse visual-text inputs in a single Transformer to reason via joint contextual encoding or enhanced cross-modal pretraining. PixelBERT (Huang et al., 2020) employs a CNN and Transformer architecture to process pixels for fine-grained reasoning (e.g., NLVR2). UniVL (Luo et al., 2020) unifies video-language reasoning with a single Transformer for temporal cross-modal tasks (e.g., TVQA). Oscar (Li et al., 2020), VinVL (Zhang et al., 2021a) anchor reasoning with object tags or enhanced visual features in a unified Transformer, boosting semantic inference (e.g., VCR, GQA). ERNIE-ViL (Yu et al., 2021) integrates scene graph knowledge into a single Transformer, enhancing compositional reasoning through structured visual-language interactions. UniT (Hu & Singh, 2021) streamlines multimodal tasks with a shared self-attention backbone for unified reasoning. PaLI (Chen et al., 2022b) scales single-Transformer reasoning with a multilingual framework for cross-lingual inference (e.g., OK-VQA). Flamingo (Alayrac et al., 2022) employs cross-attention to prioritize dynamic vision-text interactions. BEiT-3 (Wang et al., 2022b) adopts masked data modeling to unify vision-language learning. OFA (Wang et al., 2022a), BLIP-2 (Li et al., 2023c) introduce a unified multimodal framework or a querying Transformer to excel in cross-modal reasoning with improved efficiency (e.g., VQA-v2). Kosmos-1 (Huang et al., 2023b), Kosmos-2 (Peng et al., 2023) enable interleaved input processing or grounding capability for flexible multimodal understanding and precise object localization.
Multimodal LLMs-based Implicit Reasoning. This approach projects visual inputs into a large language model’s text space, leveraging the contextual reasoning capabilities of large language models (Li et al., 2023d) to improve the performance of multimodal reasoning. Their architecture contains pretrained visual encoders and large language models, arr. Vision-Encoder-LLM. CLIP-Cap (Mokady et al., 2021) projects CLIP visual features into an LLM for reasoning and captioning tasks. LLaVA (Liu et al., 2023a) enables conversational reasoning by tuning ViT-LLM integration for interactive tasks or scaling for complex VQA tasks. MiniGPT-4 (Zhu et al., 2023), InstructBLIP (Dai et al., 2023) align a ViT to a frozen LLM via a projection layer or instruction tuning, streamlining visual-text reasoning. Qwen-VL (Bai et al., 2023) incorporates a spatial-aware ViT, enhancing grounded reasoning for spatially complex tasks. mPLUG-Owl (Ye et al., 2023), LMEye (Li et al., 2024l), and Otter (Li et al., 2023a) integrate a modular visual encoder with an LLM for instruction-following and in-context learning for multimodal reasoning.
@@ 3.2.1 Prompt-based MCoT
Prompt-based Multimodal Chain-of-Thought (MCoT) methods extend the textual CoT paradigm to multimodal contexts. These methods enable step-by-step reasoning across modalities, providing strong interpretability while requiring minimal additional training. In visual reasoning, IPVR (Chen et al., 2023c) proposed a structured “see-think-confirm” prompting framework that guides LLMs through visual grounding and rationale verification. VIC (Zheng et al., 2024c) prompts textual reasoning chains before visual input to mitigate hallucinations and improve accuracy.
@@ 3.2.2 Structural Reasoning
Unlike prompt-based MCoT methods, which induce reasoning behaviour through handcrafted exemplars or a zero-shot prompting approach, structural reasoning focuses on learning reasoning patterns through supervised training. By integrating explicit procedural structures, these methods convert loosely guided reasoning into standardized, stage-wise processes. This standardization improves scalability, reliability, and efficiency in complex multimodal tasks. We categorize structural reasoning into three representative types: (i) rationale construction, which learns to produce atomic reasoning steps as interpretable scaffolds; (ii) defined reasoning procedures, which adapt structured texture reasoning schemes to multimodal settings; and (iii) modality-specific structural reasoning, which further incorporates modality-aware constraints and designs to better align with the characteristics of visual, auditory, or embodied inputs.
@@
Effective rationale learning approaches form the foundation of structural reasoning in multimodal contexts. Recent studies continue to actively explore rationale generation from various perspectives. Multimodal-CoT (Zhang et al., 2023g) proposes a two-stage Multimodal-CoT framework that decouples rationale generation from answer prediction to reduce hallucinations. T-sciq (Wang et al., 2024e) leverages teacher LLMs to generate rationales with varying complexity, showing rationale quality is key to reasoning accuracy. In autonomous driving, G-CoT (Ma et al., 2024) designs Dolphins explicitly linking rationales to visual and historical driving signals for more grounded reasoning. MC-CoT (Tan et al., 2024a) uses a self-consistency strategy to select the most accurate rationale among multiple candidates, boosting smaller models’ performance. CLoT (Zhong et al., 2024a) promotes non-linear, explorative rationale construction via Leap-of-Thought to support creative reasoning.
@@ 3.2.3 Externally Augmented Reasoning
Externally augmented reasoning introduces advanced algorithms, auxiliary tools, or expert modules to compensate for limitations in the model’s inherent reasoning capacity. These components are integrated at inference time or coupled during training, enabling more flexible, scalable, and task-specialized reasoning workflows. By decoupling core reasoning steps from the base model, such methods support long-horizon planning, precise grounding, and access to dynamic or domain-specific information. We group externally augmented methods into four categories: (i) search algorithm-enhanced MCoT, which navigates reasoning spaces via various search algorithm; (ii) tool-based augmentation, which leverages external language tools or systems to guide reasoning execution; (iii) retrieval-augmented reasoning, which incorporates relevant multimodal knowledge from external sources into the reasoning path; and (iv) multimodal enhancing, which incorporate specialized multimodal modules to support perception-driven reasoning.
@@
AGoT  (2024d) T,I Text-Image Retrieval, VQA Search Algorithm prompt aggregation and prompt flow operations Builds a graph to aggregate multi-faceted reasoning with visuals.
@@
Search strategy-driven MCoT approaches empower models to navigate and optimise reasoning trajectories throughout the reasoning process dynamically. MM-ToT (Gomez, 2023), for instance, leverages GPT-4 and Stable Diffusion, employing depth-first search (DFS) and breadth-first search (BFS) algorithms to identify the most optimal multimodal outputs according to a 0.0–1.0 metric scale. HoT (Yao et al., 2023a) creates interconnected thoughts from multimodal inputs and packages them into a single hyperedge. Unlike this, Aggregation Graph-of-Thought (AGoT) (Yang et al., 2024d) builds a reasoning aggregation graph, which integrates diverse reasoning elements at every step and subsequently incorporates visual data. Blueprint Debate on Graph (BDoG) (Zheng et al., 2024b) takes a distinctive route, discarding search algorithms and instead utilizing three agents—an affirmative debater, a negative debater, and a moderator. These agents engage in iterative debates to address multimodal questions, with the moderator ultimately synthesizing a final answer, thus implicitly constructing a graph-of-thought that explores and aggregates a wide range of thoughts. Overall, compared to prompt-based methods that rely on linear, example-driven inference, search strategy-oriented MCoT variants enable models to explore multiple reasoning pathways, thereby significantly enhancing adaptability and the depth of problem-solving.
@@ 3.3.1 Cross-Modal Reasoning
Cross-modal reasoning refers to the ability to integrate and reason across multiple modalities, such as text, images, and videos. Recent advancements in cross-modal reasoning emphasize augmenting multimodal information beyond textual inputs. These methods utilize either model-intrinsic capabilities or external tools and algorithms to achieve this augmentation. These methods aim to enhance reasoning accuracy and robustness by dynamically incorporating complementary information from diverse modalities.
@@
Beyond the use of external tools for multimodal understanding described in §3.2.3, recent approaches increasingly explore tool integration to directly facilitate multimodal reasoning. VisProg (Gupta & Kembhavi, 2023) and ProViQ (Choudhury et al., 2024) leverage program generation and procedural execution to enable cross-modal reasoning, dynamically generating executable code or logic paths to solve complex tasks such as video question answering, multi-step visual reasoning, and geometric problem-solving. In parallel, methods such as AssistGPT (Gao et al., 2023), MM-ReAct (Yang et al., 2023), and Multi-Modal-Thought (Lin et al., 2025a) adopt modular integration frameworks (e.g., PEIL, vision expert prompting) to coordinate tool use based on reasoning progression. These systems enable interpretable and adaptive reasoning by calling different tools dynamically during task execution. VisualReasoner (Cheng et al., 2024a) further introduces a data synthesis strategy to generate multi-step reasoning traces, which are then used to train a plug-and-play visual reasoning module applicable to a variety of vision-language backbones. Collectively, these efforts extend the landscape of multimodal reasoning by combining program induction, dynamic tool orchestration, and data-driven reasoning supervision.
@@
Besides external tools, algorithmic methods have also been explored to enhance cross-modal reasoning through explicit cognitive process modelling. FAST (Sun et al., 2024a) and ICoT (Gao et al., 2024a) leverage cognitive processes akin to human thinking. Specifically, FAST employs a system switch adapter to dynamically alternate between fast and slow thinking modes, whereas ICoT utilizes Attention-driven Selection (ADS) to interleave visual and textual reasoning steps. Meanwhile, Image-of-Thought (Zhou et al., 2024b) and CoTDiffusion (Ni et al., 2024a) focus on generating visual rationales. Image-of-Thought extracts visual information step-by-step, whereas CoTDiffusion creates visual subgoal plans, extending algorithmic augmentation into robotics domains.
@@
In contrast to approaches relying on external augmentation, several methods exploit intrinsic model capabilities to achieve cross-modal reasoning. These approaches rely on the inherent capabilities of LMMs to generate or infer multimodal information without relying on external tools. T-SciQ (Wang et al., 2024e), Visual-CoT (Rose et al., 2023) and VoCoT (Li et al., 2024m) demonstrated how fine-tuning LMMs on carefully designed CoT datasets (e.g., VoCoT-Instruct80K) could enable single-step multimodal reasoning in charts, documents, and geometry problems. MVoT (Li et al., 2025b) represents an early effort, where a self-contained architecture iteratively refines visual-textual representations for embodied reasoning tasks.
@@
Table 5: Approaches of Multimodal-o1. It mainly relies on a multi-stage, structured reasoning path to solve problems.
@@
RBF++ (2025d) LLaMA3-8B-Instruct GSM8K + SVAMP + MATH T SR-MCTS + PPRM Math Proposes SR-MCTS for structured search and PPRM for evaluating reasoning boundaries